OpenStudy (anonymous):

The 6 consecutive integers below add up to 447. x-2 x-1 x x+1 x+2 x+3 What is the value of x?

OpenStudy (anonymous):

divide 447 by 6 subtract .5 do you see why?

OpenStudy (anonymous):

I see why you would divide by 6 but not subtract by 5

OpenStudy (anonymous):

subtract 0.5

OpenStudy (anonymous):

if you add 6 consecutive numbers together, the average will be right between the third and fourth

OpenStudy (ajprincess):

x-2+x-1+x+x+1+x+2+x+3=447 6x+3=447 solve for x. Does that help?

OpenStudy (anonymous):

How did you get 6x+3

OpenStudy (anonymous):

she added up the 6 consecutive integers in terms of x and set it equal to 447

OpenStudy (anonymous):

I see now, the answer is 74

OpenStudy (ajprincess):

the 6 consecutive integers interms of x are given to u. nd also it is given that their sum is 447. Adding them and equating to 447 gives u 6x+3=447

OpenStudy (ajprincess):

yup that's right:)

OpenStudy (anonymous):

ajprincess's way is the sure way, my way is the quick way.

OpenStudy (anonymous):

thanks

OpenStudy (ajprincess):

welcome:)